OpenStudy (anonymous):

HELP!! the expression 6y^2 + 13y-5 reprsents the area of a parallelgram. Which xpression could represent the length of the base the parallelogram?

OpenStudy (cwrw238):

6y^2 + 13y-5 = (3y - 1)(2y + 5) - one of these factors could be the base of the parallelogram the other would be its perpendicular height

OpenStudy (anonymous):

but i still dont understand how to find the length

OpenStudy (cwrw238):

the length of base could be 3y - 1 or 2y + 5

OpenStudy (anonymous):

can you explain to me hoe you got that answer

OpenStudy (anonymous):

how*

OpenStudy (cwrw238):

area of parallelogram is given as 6y^2 + 13y-5 and this area = length base * perpendicular height since 6y^2 + 13y-5 = (3y - 1) times (2y + 5) either one could be length of base.

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

yw